HUBBLE SPACE TELESCOPE DAILY REPORT #1882 PERIOD COVERED: 0000Z (UTC) 05/20/97 - 0000Z (UTC) 05/21/97 Daily Status Report as of 141/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of NIC/1 7152 (NICMOS Absolute Photometry Test) The Near Infrared Camera (Camera One) was used to conduct an absolute photometric calibration test. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.2 Completed STIS/CCD 7076 (STIS Corrector Alignment, Fine) The Space Telescope Infrared Spectrograph (CCD) was used to make observations of BPM16274 as part of a fine correction alignment and calibration procedure. This is to refine the STIS corrector system alignment so that the image quality at the spectrographs slit plane is optimized. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/MA1/MA2 7084 (Verification of Optical Format of STIS MAMA Modes) The Space Telescope Infrared Spectrograph (MA1 and MA2) was used to perform verification calibration observations of the optical format for the MAMA modes and extreme MSM positions using the Cr/Pt/Ne lamp. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6362 (Supernova Host Galaxy) The WF/PC-2 was used to make observations of the supernova host galaxy GAL-160828. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/1/2/3 7043 (NICMOS Focus Monitor) The Near Infrared Cameras (Cameras One, Two and Three) were used to make observations of NGC-3603 in order to make focus calibration and monitoring measurements. This proposal is built to run NIC/1/2 focus sweeps on a weekly basis, and NIC /3 focus sweeps twice a week during SMOV. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/MA2 7093 (Dark Measurement for MAMA Band 1, and Band 2) The Space Telescope Infrared Spectrograph (MA2) was used to make dark calibration observations of the MAMA Band 1 and Band 2. These observations are to be used to establish that the detectors are performing nominally, that no internal movement of the MCP has taken place, and to establish the dark level in orbit for a given operating temperature. The observations were completed as planned, and no anomalies were reported. 1.7 Completed Four Sets of WF/PC-2 7553 (Pure Parallel Archive Survey II) WF/PC-2 parallel observations were obtained in the ultraviolet in search for interesting objects. This program is designed to find extragalactic sources with near ultraviolet excesses (3000 A0) outside the galactic plane and far ultraviolet sources (1600 A0) in the galactic plane. The procedure was executed as scheduled, and no anomalies were reported. 1.8 Completed Two Sets of WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats) The WF/PC-2 was used to make observations of the bright Earth to obtain streak flats which are used to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/2 7154 (NICMOS Coronographic Hole Monitor) The Near Infrared Camera (NIC Two) was used to make observations of object NGC-3603 in order to monitor the effects of coronagrapghic hole. The purpose of this proposal is to take lamp flats with camera Two (and cameras One and Three occasionally) to monitor the location of the coronographic spot. The observations were completed as scheduled, and no anomalies were reported. 1.10 Completed FOC/96 6834 (Lenses, Mirrors, and HST: IRAS FSC10214+4724 Under a Compound Microscope) The Faint Object Camera (f/96) was used to make observations of IRAS FSC10214+4724. Our existing WF/PC-2 images using filter F814W provides compelling evidence that this object which is at redshift of 2.3 and is gravitationally lensed. The observations were completed as planned, and no anomalies were reported. 1.11 Completed Two sets of WF/PC-2 6610 (WF/PC-2 Lyman-Alpha Imaging of Galaxy cluster at Z=2.4) The WF/PC-2 was used to make observations of an apparent blank field of view hoping to make Lyman-Alpha observations of faint galaxies located at a distance of Z=2.4. Earlier exposures revealed many compact Lyman-Alpha emitters at this distance. These observations were executed in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 14 Successful : 14 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Operations Notes: The MAMA2 high voltage was successfully turned on at 141/0255Z, but some concerns, detailed in HSTAR 6074, were noted. Data from other turn-ons indicate a possible correlation between the MAMA2 tube temperature and the additional noise event counts resulting from the glass phosphorescence. During the initial high voltage ramp, the temperature was around 21-22 degC, resulting in noise event counts of around 2000 per second. As the tube temperature continued to increase to about 31 degC, the noise event counts increased significantly. The high voltage ramped down nominally with the low voltages now being managed via the SMS so they are always on external to the SAAs. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH